# 3D-RAD: A Comprehensive 3D Radiology Med-VQA Dataset with Multi-Temporal Analysis and Diverse Diagnostic Tasks

### v1

@@ 5 Conclusion
Appendix A Technical Appendices and Supplementary Material
Our dataset and code are publicly available at https://github.com/Tang-xiaoxiao/M3D-RAD.
A.1 Limitations and Future Work.
@@ Appendix A Technical Appendices and Supplementary Material
A.2 LLM Scoring Consistency Evaluation.
We assess the consistency of scoring across four large language models (LLMs). The left panel in  Figure 8,Figure 8 shows the average pairwise agreement scores for each model relative to others, while the right panel presents their rankings based on overall agreement. Notably, GPT-4o-mini achieves the highest consistency, aligning closely with other models, particularly in the high-score range. Based on this result, we adopt GPT-4o as the default evaluator in our scoring pipeline.
@@
A.3 The Importance of Tasks 5-6
A.4 Question Construction Prompts and Templates.
Figure 9,Figure 10,Figure 11,Figure 12,Figure 13,Figure 14 illustrate the complete set of carefully designed templates used for question construction across different tasks. Each template is tailored to the unique characteristics and objectives of the corresponding task, enabling accurate and diverse generation of clinically meaningful question-answer pairs.
@@ A.3 The Importance of Tasks 5-6
A.5 LLM-Based Scoring Prompt.
@@
A.6 Finetuned Results of M3D-4B Models.
@@
Task1: Anomaly Detection BLEU 16.31 20.54 29.35 39.66
@@
Task2: Image Observation BLEU 15.06 21.10 30.54 33.28
Rouge 23.19 26.03 37.53 42.45
Meteor 16.56 21.18 29.12 31.38
BERTScore 87.11 88.13 89.78 90.72
Task3: Medical Computation BLEU 2.55 5.65 15.23 20.85
Rouge 5.63 6.91 23.76 30.38
Meteor 4.20 4.53 14.06 16.92
BERTScore 85.74 84.89 90.70 93.25
@@
Task5: Static Temporal Diagnosis Accuracy 24.31 61.01 74.19 74.77
Task6: Longitudinal Temporal Diagnosis Accuracy 25.40 41.17 48.11 49.30
A.7 Case Visualization.
Figure 16: An example of QA Construction
A.8 Zero-shot Subtask Performance.
Figure 17,Figure 18,Figure 19,Figure 20,Figure 21 illustrates the performance of models under the zero-shot setting across different subtasks. The results reveal substantial variation in accuracy, highlighting that some subtasks are more tractable in the absence of fine-tuning, while others remain highly challenging. This emphasizes the varying complexity and reasoning demands posed by each subtask within our benchmark.
@@
A.9 Finetuned Subtask Performance.
Figure Figure 22,Figure 23,Figure 24,Figure 25,Figure 26 illustrate the performance comparison across subtasks after fine-tuning. The results highlight consistent improvements in most subtasks, demonstrating the effectiveness of our domain-specific training data in enhancing the model’s task-specific capabilities.Notably, the most significant gains are observed in subtasks requiring temporal reasoning, suggesting that our dataset effectively supports learning nuanced clinical progressions.
Figure 22: Finetuned BLEU Result of Task1-3
Figure 23: Finetuned F1 Result of Task1-3
Figure 24: Finetuned results of models on Task 4
Figure 25: Finetuned Result of Task4
Figure 26: Finetuned Result of Task5-6
A.10 Word Cloud Visualizations Across Tasks.
To illustrate the diversity of our benchmark, we present word cloud visualizations for each individual task. These visualizations highlight the varying distributions of clinical concepts across tasks, reflecting the distinct focuses and linguistic patterns inherent to each QA setting. Figure Figure 27 shows the task-wise word clouds.
Task 1
Task 2
Task 3
Task 4
Task 5
Task 6
A.11 Detailed Analysis of Experimental Results.
In the main text, we present a summary of the overall experimental results (Table 3, Table 4). Here, we provide a more detailed analysis of these results.
A.11.1 Result Analysis
@@
A.11.2 Failure Case Analysis
@@
A.12 Model-Level Scoring Consistency Analysis.
A.12.1 Model Selection
A.12.2 Evaluation Metrics
A.12.3 Evaluation Procedure
We then perform pairwise comparisons across all models using the metrics defined in Section A.11.2. For each model, we compute its average metric value against the other three models.  Figure 8 illustrates the average metric values of each model relative to the others. Figure Figure 8 shows the ranking of each model based on these averaged metrics.
A.13 Human Annotation Protocol.
A.13.1 Annotator Background
Figure 29: Guideline
A.13.2 Annotation Procedure
A.14 Original Dataset and QA Construction Case Visualization.
A.14.1 Original Dataset
Based on these sources:
A.14.2 QA Construction Case Visualization

### v2

@@ 5 Conclusion
6 Acknowledgments
This work is supported by the National Key R&D Program of China (Grant No. 2024YFC3308304), the "Pioneer" and "Leading Goose" R&D Program of Zhejiang (Grant no. 2025C01128), the National Natural Science Foundation of China (Grant No. 62476241), the Natural Science Foundation of Zhejiang Province, China (Grant No. LZ23F020008).
Appendix A Limitations and Future Work.
@@ 6 Acknowledgments
Appendix B LLM Scoring Consistency Evaluation Details.
We assess the consistency of scoring across four large language models (LLMs). The left panel in  Figure 8, and Figure 8 shows the average pairwise agreement scores for each model relative to others, while the right panel presents their rankings based on overall agreement. Notably, GPT-4o-mini achieves the highest consistency, aligning closely with other models, particularly in the high-score range. Based on this result, we adopt GPT-4o as the default evaluator in our scoring pipeline.
@@
B.1 Model Selection
B.2 Evaluation Metrics
Appendix C Evaluation Procedure.
We then perform pairwise comparisons across all models using the metrics defined in Section B.2 For each model, we compute its average metric value against the other three models. Figure 8 illustrates the average metric values of each model relative to the others. Figure 8 shows the ranking of each model based on these averaged metrics.
Appendix D Human Annotation Protocol.
D.1 Annotator Background
D.2 Annotation Procedure
Table 5: Results across General VLMs on 3D-RAD Benchmark.
Model Task1 (ROUGE1 / BERTScore) Task2 (ROUGE1 / BERTScore) Task3 (ROUGE1 / BERTScore) Task4 (ACC) Task5 (ACC) Task6 (ACC)
llava-onevision(7b) 22.34 / 86.24 26.30 / 86.79 5.93 / 91.81 29.67 6.86 6.86
qwen2.5-vl(3B) 24.40 / 84.96 22.37 / 83.77 18.21 / 92.31 22.75 24.09 24.09
qwen2.5-vl(32B) 21.77 / 87.66 20.30 / 87.64 11.46 / 90.73 58.60 28.62 28.36
gemma-3(4b) 22.91 / 87.44 22.50 / 87.41 8.56 / 90.50 21.47 10.75 10.74
gemma-3(27b) 26.21 / 88.15 30.43 / 88.89 10.55 / 90.54 28.57 24.09 24.09
gemini-2.0-flash 24.70 / 88.93 27.93 / 89.29 0.53 / 84.36 40.42 23.08 21.33
intern-vl-3(8B) 28.44 / 88.83 34.31 / 89.63 10.65 / 91.64 55.30 26.63 26.61
gpt-4.1-nano 10.99 / 85.00 11.41 / 85.32 4.31 / 83.63 37.95 19.95 25.11
gpt-4.1 15.08 / 86.03 16.55 / 86.50 2.76 / 82.82 61.98 34.08 68.05
Ours-Finetuned (Llama2-7B) 33.76 / 89.16 39.12 / 90.00 36.06 / 94.65 81.09 51.20 74.78
Ours-Finetuned (Phi3-4B) 42.45 / 90.72 50.52 / 92.19 36.46 / 94.86 82.43 49.30 74.77
Figure 9: An example of QA Construction
Appendix E Prompts and Templates.
E.1 Question Construction Prompts and Templates.
Figure 11, Figure 12, Figure 13, Figure 14, Figure 15, and Figure 16 illustrate the complete set of carefully designed templates used for question construction across different tasks. Each template is tailored to the unique characteristics and objectives of the corresponding task, enabling accurate and diverse generation of clinically meaningful question-answer pairs.
Figure 10: Guideline
@@ D.2 Annotation Procedure
E.2 LLM-Based Scoring Prompt.
@@
Appendix F Original Dataset.
Based on these sources:
Appendix G QA Construction Case Visualization.
Appendix H Detailed Analysis of Experimental Results.
In the main text, we present a summary of the overall experimental results (Table 3 and Table 4). Here, we provide a more detailed analysis of these results.
H.1 Finetuned Results of M3D-4B Models.
@@
Task1: Anomaly Detection BLEU 15.06 21.10 30.54 33.28
Rouge 23.19 26.03 37.53 42.45
BERTScore 87.11 88.13 89.78 90.72
Task2: Image Observation BLEU 16.31 20.54 29.35 39.66
@@
Task3: Medical Computation BLEU 2.55 7.01 25.47 33.52
Rouge 5.63 9.95 31.84 36.46
BERTScore 85.74 85.97 93.06 94.86
@@
Task5: Static Temporal Diagnosis Accuracy 25.40 41.17 48.11 49.30
Task6: Longitudinal Temporal Diagnosis Accuracy 24.31 61.01 74.19 74.77
H.2 Finetuned Subtask Performance.
Figure 18, Figure 19, Figure 20, Figure 21, and Figure 22 illustrate the performance comparison across subtasks after fine-tuning. The results highlight consistent improvements in most subtasks, demonstrating the effectiveness of our domain-specific training data in enhancing the model’s task-specific capabilities.Notably, the most significant gains are observed in subtasks requiring temporal reasoning, suggesting that our dataset effectively supports learning nuanced clinical progressions.
Figure 18: Finetuned BLEU Result of Task 1-3
Figure 19: Finetuned F1 Result of Task 1-3
Figure 20: Finetuned results of models on Task 4
Figure 21: Finetuned Result of Task 4
Figure 22: Finetuned Result of Task 5-6
H.3 Zero-shot Subtask Performance.
Figure 23, Figure 24, Figure 25, Figure 26, and Figure 27 illustrates the performance of models under the zero-shot setting across different subtasks. The results reveal substantial variation in accuracy, highlighting that some subtasks are more tractable in the absence of fine-tuning, while others remain highly challenging. This emphasizes the varying complexity and reasoning demands posed by each subtask within our benchmark.
@@
H.4 Result Analysis
@@
Appendix I Failure Case Analysis
@@
Appendix J Word Cloud Visualizations across Tasks.
To illustrate the diversity of our benchmark, we present word cloud visualizations for each individual task. These visualizations highlight the varying distributions of clinical concepts across tasks, reflecting the distinct focuses and linguistic patterns inherent to each QA setting. Figure 29 shows the task-wise word clouds.
Appendix K Additional Evaluations on General VLMs
We also evaluate our benchmark on mainstream general-purpose vision–language models.As illustrated in Table 5, though some models perform better—especially on tasks 4–6 with more reliable metrics, none achieve strong results across all tasks. Models fine-tuned on 3D-RAD consistently outperform general models, highlighting the dataset’s value in promoting domain-specific learning.
Table 7: Evaluation accuracy on Task4–Task6 under different finetuning settings.
Finetuned Task(s) Evaluate Task4 (ACC) Evaluate Task5 (ACC) Evaluate Task6 (ACC)
Only Task1 82.03 38.78 72.61
Only Task2 81.88 48.73 73.55
Only Task3 81.94 48.30 72.78
Only Task4 81.96 38.97 73.52
Zero-Shot M3D (Llama2-7B) 18.00 25.47 24.17
Zero-Shot M3D (Phi3-4B) 40.25 25.40 24.31
All Tasks (Ours) 82.43 49.30 74.77
Table 8: Evaluation metrics on Task1 (BLEU / ROUGE1 / F1) under different finetuning settings.
Finetuned Task(s) Evaluate Task1 (BLEU / ROUGE1 / F1)
Only Task5 31.60 / 42.01 / 90.48
Only Task6 32.45 / 42.87 / 90.73
Zero-Shot M3D (Llama2-7B) 9.10 / 18.64 / 86.07
Zero-Shot M3D (Phi3-4B) 15.06 / 23.19 / 87.11
All Tasks (Ours) 33.28 / 42.45 / 90.72
Appendix L Task-level Ablation Studies
We added task-level ablation studies to examine cross-task effectiveness in Table 7 and Table 8. The table below presents results where each task is fine-tuned individually and then evaluated on all other tasks. As shown, fine-tuning on a single task can still improve performance on other tasks. However, fine-tuning on the full dataset consistently yields the best results. This demonstrates that the effectiveness of our dataset stems not only from its size but also from the inclusion of high-quality domain knowledge.
Task 1
Task 2
Task 3
Task 4
Task 5
Task 6